# VidGuard-R1: AI-Generated Video Detection and Explanation via Reasoning MLLMs and RL

### v2

@@ Abstract
With the rapid advancement of AI-generated videos, there is an urgent need for effective detection tools to mitigate societal risks such as misinformation and reputational harm. In addition to accurate classification, it is essential that detection models provide interpretable explanations to ensure transparency for regulators and end users. To address these challenges, we introduce VidGuard-R1, the first video authenticity detector that fine-tunes a multi-modal large language model (MLLM) using group relative policy optimization (GRPO). Our model delivers both highly accurate judgments and insightful reasoning. We curate a challenging dataset of 140k real and AI-generated videos produced by state-of-the-art generation models, carefully designing the generation process to maximize discrimination difficulty. We then fine-tune Qwen-VL using GRPO with two specialized reward models that target temporal artifacts and generation complexity. Extensive experiments demonstrate that VidGuard-R1 achieves state-of-the-art zero-shot performance on existing benchmarks, with additional training pushing accuracy above 95%. Case studies further show that VidGuard-R1 produces precise and interpretable rationales behind its predictions. The code is publicly available at https://VidGuard-R1.github.io.
@@ 1 Introduction
In the past year, we have witnessed unprecedented progress in video generation models, with dramatic improvements in realism and quality. The release of powerful models such as Sora (Brooks et al., 2024), Wan (Wang et al., 2025a), and HunyuanVideo (Kong et al., 2024) has made AI-generated videos more accessible to the public, further blurring the line between synthetic videos and real ones. At the same time, these advancements have led to a series of social risks, including the spread of misinformation, violations of privacy rights, damage to personal reputations, and increased susceptibility to scams and fraud.
Motivated by its practical significance, several pioneering works have been developed to detect AI-generated videos. Early approaches primarily targeted DeepFake-style facial forgeries (Qian et al., 2020b; Tan et al., 2024; Gu et al., 2021), which often assumed single-subject, frontal-face scenarios under constrained settings. These assumptions diverge significantly from open-domain, multi-scene videos produced by modern generative models. More recent detectors leverage spatial-temporal consistency (Ma et al., 2024; Bai et al., 2024b; Liu et al., 2024); however, such methods are limited in capturing higher-level semantic or causal inconsistencies and can be easily bypassed by post-processing techniques. Other methods are trained on curated fake video detection datasets (Chen et al., 2024a; Ni et al., 2025; Kundu et al., 2025), but these benchmarks often lack coverage of newly emerging models and fail to reflect the full diversity of generative capabilities. A recent benchmark (Chen et al., 2024a) shows that even state-of-the-art detectors still struggle to reliably identify videos from advanced models like Sora. Furthermore, these detectors typically offer only binary decisions without accompanying explanations, which raises concerns for transparency, especially when detection outcomes affect content moderation or legal accountability. Users are also more likely to trust detection systems that provide interpretable reasoning.
@@
To address this, we adopt reinforcement learning (RL), which has shown promise in enhancing LLM reasoning (Guo et al., 2025). Notably, Video-R1 (Feng et al., 2025) outperforms commercial models on video reasoning tasks. RL enables MLLMs to develop self-improving reasoning via outcome-based rewards. We hypothesize that RL fine-tuning can help models detect subtle temporal and generative artifacts. Key to this is designing effective reward models. Simple binary rewards (e.g., 1 for real, 0 for fake) are insufficient. Instead, we propose two strategies: (1) injecting temporal artifacts into both real and fake videos to encourage temporal reasoning, and (2) assigning higher rewards to videos generated with more diffusion steps, which are harder to detect. Incorporating these into a group relative policy optimization (GRPO) framework leads to over 86% accuracy on our dataset and 95% accuracy on two benchmarks.
@@ 2.1 AI-generated video detection method
In recent years, research on AI-generated video detection has primarily focused on deepfake videos featuring synthetic faces (Pei et al., 2024), with methods based on spatial-temporal consistency, frequency artifacts, and data-driven approaches. However, these approaches often struggle to generalize beyond face-centric content to more diverse, real-world videos. Recently, general video detection methods have emerged: AIGDet (Bai et al., 2024a) captures spatial-temporal anomalies, DeCoF (Ma et al., 2024) exploits frame consistency, and diffusion-based representations have been used to track temporal dynamics (Liu et al., 2024). Other works identify key factors such as appearance, motion, and geometry for classifier training (Chang et al., 2024). MLLMs have also been explored for visual forgery detection: FakeShield (Xu et al., 2024) employs supervised fine-tuning (SFT) for image forgery detection and relies on GPT-4o for explanation generation, while SafeWatch (Chen et al., 2024b) applies both SFT and direct preference optimization (DPO) to train a video guardrail model with transparent reasoning. In contrast, our work is the first to fine-tune a multi-modal large language model (MLLM) using group relative policy optimization (GRPO) for AI-generated video detection, demonstrating strong generalization across a wide range of recent video generative models and benchmark datasets.
@@ 2.2 AI-generated video detection dataset
As this is a very recent area of research, only a few benchmarks have been proposed. The generated video dataset (GVD) (Bai et al., 2024a) (11k samples) and GenVideo (Chen et al., 2024a) (with millions of samples) consider settings where both training and test videos are generated by the same series of models. However, these benchmarks lack prompt/image–video pairs, semantic labels, or cross-source settings. GVF (2.8k samples) contains prompts/images–video pairs and semantic labels, but does not provide cross-source settings. GenVidBench (Ni et al., 2025) consists of 100k videos and incorporates cross-source settings, but the video generation models used are less advanced, such as CogVideo and SVD. To address these limitations, we construct a curated dataset of 140,000 real–fake video pairs generated with state-of-the-art video generation models (HunyuanVideo (Kong et al., 2024) and CogVideoX (Yang et al., 2024)), standardizing video properties to mitigate superficial cues and encourage models to focus on intrinsic visual realism.
@@ 3.1.1 Data construction for video realism discrimination
High-quality training data is essential for video reasoning in MLLMs.. However, many existing benchmarks for real vs. generated video classification, such as GenVideo (Chen et al., 2024a) and GenVidBench  (Ni et al., 2025), exhibit uncontrolled discrepancies in basic metadata—e.g., real videos are often longer than 10 seconds while generated ones are typically under 4 seconds in GenVideo. Moreover, they reveal clear modality-level gaps in motion dynamics and content contrasts between real and generated videos. These differences introduce unintended shortcuts, enabling models to rely on superficial cues like duration or resolution rather than actual visual realism. As a result, VidGuard-R1 attains over 96% accuracy on both GenVideo and GenVidBench by exploiting such artifacts. To mitigate this reward hacking behavior, we construct a curated dataset with standardized video properties, encouraging models to focus on intrinsic visual content.
We collect real videos from the InternVid (Wang et al., 2023c) and ActivityNet (Caba Heilbron et al., 2015) datasets and generate their corresponding fake counterparts using HunyuanVideo (Kong et al., 2024) and CogVideoX (Yang et al., 2024). We specifically choose these two models because they support conditioning on both the first-frame image and a text description—an essential requirement for generating videos that are contextually aligned with their real counterparts. To achieve such alignment, we provide the generation models with the first frame of each real video along with a textual caption describing its content. For ActivityNet, which lacks native captions, we extract concise descriptions using Qwen2.5-VL 72B. This pairing strategy mitigates content-based biases and forces the model to reason over subtle visual details.
@@ 3.3.3 GRPO with quality evolutionary videos (GRPO-Q)
This reward formulation enables the model to move beyond binary discrimination and perform fine-grained analysis of video quality. By learning to associate subtle differences in generation steps with quality variations, the model develops a deeper understanding of the diffusion process and its impact on perceptual realism. As a result, it can not only detect whether a video is fake, but also infer but also estimate the degree of quality degradation in generated videos. This facilitates more interpretable and controllable evaluation of generated content quality.
@@ 4.1.3 Training Setup
We employ Qwen2.5-VL-7B as the base MLLM and conduct all experiments on four NVIDIA A100 GPUs (80GB). Each video is represented by up to 16 frames, where each frame is resized to a $28\times 28$ spatial resolution and mapped to 128 feature channels for encoder input during both training and inference. For GenVideo and GenVidBench, we follow their official evaluation protocols and adopt 8-frame inputs. In GRPO training, we sample 8 responses per input; for GRPO-TA, we additionally sample 4 responses from temporally manipulated variants of the input to enhance robustness against temporal artifacts. Training proceeds in two stages: first, the base model is fine-tuned for one epoch on the CoT dataset, yielding the SFT-CoT MLLM; second, we initialize VidGuard-R1 with SFT-CoT and perform reinforcement learning for approximately 2,000 steps.
@@
VidGuard-R1 (GRPO,
GenVideo-pretrained,
Zero-shot)
MLLM 97.24 96.59 99.88 99.93 88.14 96.37
@@ 4.2.1 Our dataset
We evaluate VidGuard-R1 on our dataset with several methods, including CNN-based models (SlowFast (Feichtenhofer et al., 2019), I3D (Carreira & Zisserman, 2017), TRN (Zhou et al., 2018)), Transformer-based models (UniFormer V2 Li et al. (2022a), TimeSformer (Bertasius et al., 2021), VideoSwin (Liu et al., 2022), MViT V2 (Li et al., 2022b)), and MLLM-based models (Qwen2.5-VL (Bai et al., 2025) and GPT-4.1 mini (OpenAI, 2025)). For CNN and Transformer models, we use the default training settings provided by the MMAction2 framework (Contributors, 2020).
@@ 4.2.2 GenVidBench benchmark
VidGuard-R1 (GRPO,
GenVidBench-pretrained, Zero-shot)
MLLM R 0.95 0.98 0.90 0.89 0.97 0.85 0.99 0.93 0.81 0.87 0.92
@@ 4.2.3 GenVideo benchmark
The GenVideo dataset comprises approximately 2.2M training samples and 20k testing samples, with generated videos sourced from a diverse set of models, including Sora (OpenAI, 2024), MorphStudio (mor, 2025), Gen2 (Esser et al., 2023b), HotShot (hot, 2025), Lavie (Wang et al., 2025b), Show-1 (Zhang et al., 2024a), MoonValley (moo, 2025), Crafter (Chen et al., 2023), ModelScope (Wang et al., 2023a), and WildScrape (Wei et al., 2024). Following the official evaluation protocol, we benchmark two MLLM baselines and three variants of VidGuard-R1. Among these, VidGuard-R1 (GRPO) consistently outperforms almost all prior detection methods across videos generated by the various models. As shown in Table 3, it achieves an F1 score improvement of 0.06 compared to DeMamba-XCLIP. Complete detection model results are provided in Appendix B.
@@ 4.2.5 Ablation study
Method
Expl. Score
Expl. Score
@@
Explanation quality and accuracy comparison
Table 4 presents results on the HunyuanVideo (Kong et al., 2024) and CogVideoX (Yang et al., 2024) datasets. We report explanation quality scores, which are rated on a 1–10 scale (with 10 indicating excellent quality and full alie code is publicly availagnment) by GPT-4.1 mini using the LLM-as-a-Judge prompt described in Appendix C. Compared to baseline models such as Qwen2.5-VL-7B and GPT-4.1 mini, our VidGuard-R1 GRPO variants achieve consistent improvements in both classification accuracy and explanation quality.
@@
GRPO-TA reward ablation
@@
GRPO-Q reward ablation
Table 6 presents an ablation study on GRPO-Q conducted on our dataset by varying the number of intermediate diffusion steps included per real video during fine-tuning. Using more steps provides richer supervision of video quality progression, improving detection accuracy. The best accuracy of 84.05% is obtained with five steps, which is the setting used in our main experiments.

### v3

@@ Abstract
The rapid proliferation of AI-generated video necessitates robust detection tools that offer both high accuracy and human-interpretable explanations. While existing MLLM-based detectors rely on supervised fine-tuning (SFT) or direct preference optimization (DPO), these methods are often bottlenecked by static, pre-labeled datasets that fail to capture the evolving, multi-step physical inconsistencies of modern generative models. To bridge this gap, we introduce VidGuard-R1, the first video authenticity detector to utilize group relative policy optimization (GRPO). Moving beyond passive preference matching, VidGuard-R1 employs a reinforcement learning framework that encourages the model to explore and rank multiple reasoning paths. By introducing specialized reward models for temporal stability and diffusion-aware complexity, we incentivize the model to discover ’physics-grounded’ artifacts. Our contributions include: (1) a curated dataset of 140,000 challenging real/fake video pairs; (2) a GRPO-based training paradigm that achieves state-of-the-art zero-shot performance; and (3) a reasoning-first architecture that provides precise, verifiable rationales for its forensic judgments. Project website: https://vidguard-r1.github.io/
@@ 1 Introduction
In the past year, we have witnessed unprecedented progress in video generation models, with dramatic improvements in realism and quality. The release of powerful models such as Sora (Brooks et al., 2024), Wan (48), and HunyuanVideo (Kong et al., 2024) has made AI-generated videos more accessible to the public, further blurring the line between synthetic videos and real ones. At the same time, these advancements have led to a series of social risks, including the spread of misinformation, violations of privacy rights, damage to personal reputations, and increased susceptibility to scams and fraud.
Motivated by its practical significance, several pioneering works have been developed to detect AI-generated videos. Early approaches primarily targeted DeepFake-style facial forgeries (Qian et al., 2020; Tan et al., 2024; Gu et al., 2021), which often assumed single-subject, frontal-face scenarios under constrained settings. These assumptions diverge significantly from open-domain, multi-scene videos produced by modern generative models. More recent detectors leverage spatial-temporal consistency (Ma et al., 2024; Bai et al., 2024; Liu et al., 2024); however, such methods are limited in capturing higher-level semantic or causal inconsistencies and can be easily bypassed by post-processing techniques. Other methods are trained on curated fake video detection datasets (Chen et al., 2024a; Ni et al., 2025; Kundu et al., 2025), but these benchmarks often lack coverage of newly emerging models and fail to reflect the full diversity of generative capabilities. A recent benchmark (Chen et al., 2024a) shows that even state-of-the-art detectors still struggle to reliably identify videos from advanced models like Sora. Furthermore, these detectors typically offer only binary decisions without accompanying explanations, which raises concerns for transparency, especially when detection outcomes affect content moderation or legal accountability. Users are also more likely to trust detection systems that provide interpretable reasoning.
@@
To address this, we adopt reinforcement learning (RL), which has shown promise in enhancing LLM reasoning (Guo et al., 2025). Notably, Video-R1 (Feng et al., 2025) outperforms commercial models on video reasoning tasks. RL enables MLLMs to develop self-improving reasoning via outcome-based rewards. We hypothesize that RL fine-tuning can help models detect subtle temporal and generative artifacts. Key to this is designing effective reward models. Simple binary rewards (e.g., 1 for real, 0 for fake) are insufficient. Instead, we propose two strategies: (1) injecting temporal artifacts into both real and fake videos to encourage temporal reasoning, and (2) assigning higher rewards to videos generated with more diffusion steps, which are harder to detect. Incorporating these into a group relative policy optimization (GRPO) framework yields over 86% accuracy on our dataset and 95% on two benchmarks.
@@ 2.1 AI-generated video detection method
Recent research on AI-generated video detection has largely focused on deepfake videos with synthetic faces (Pei et al., 2024), using spatial–temporal consistency, frequency artifacts, or data-driven approaches. These methods often struggle to generalize beyond face-centric content to more diverse, real-world videos. Recently, general video detection methods have emerged: AIGDet (Bai et al., 2024) captures spatial–temporal anomalies, DeCoF (Ma et al., 2024) exploits frame consistency, and diffusion-based representations track temporal dynamics (Liu et al., 2024). Other works identify appearance, motion, and geometry as key factors for classifier training (Chang et al., 2024).
Recent efforts in forgery detection leverage multimodal LLMs, including FakeShield (Xu et al., 2024), which employs supervised fine-tuning, and SafeWatch (Chen et al., 2024b), which applies direct preference optimization (DPO) for video guardrails. However, DPO relies on static preference pairs that struggle to capture subtle temporal inconsistencies in evolving generative models. We introduce the first application of group relative policy optimization (GRPO) to AI-generated video detection. By enabling iterative exploration and ranking of multiple reasoning paths, our method promotes a deeper physics-aware understanding of video consistency, leading to improved generalization across frontier generative models and diverse benchmarks.
@@ 2.2 AI-generated video detection dataset
Given the recency of this research area, only a limited number of benchmarks have been introduced. The generated video dataset (GVD) (Bai et al., 2024) (11k samples) and GenVideo (Chen et al., 2024a) (with millions of samples) consider settings where both training and test videos are generated by the same series of models. However, these benchmarks lack prompt/image–video pairs, semantic labels, or cross-source settings. GVF (2.8k samples) contains prompts/images–video pairs and semantic labels, but does not provide cross-source settings. GenVidBench (Ni et al., 2025) consists of 100k videos and incorporates cross-source settings, but the video generation models used are less advanced, such as CogVideo and SVD.
Moreover, existing datasets often contain shortcuts in resolution, frame rate, bitrate, or data-source imbalance, enabling models to exploit superficial cues rather than learn intrinsic visual realism. To address these limitations, we construct a curated dataset of 140,000 real–fake videos generated with state-of-the-art video generation models: HunyuanVideo (Kong et al., 2024) and CogVideoX (Yang et al., 2024). Our dataset explicitly standardizes bitrate, resolution, frame rate, and content distribution, resulting in a shortcut-free benchmark that encourages models to rely on semantic and temporal realism rather than on metadata artifacts.
@@ 3.1.1 Data construction for video realism discrimination
High-quality training data is essential for video reasoning in MLLMs. However, many existing benchmarks for real vs. generated video classification, such as GenVideo (Chen et al., 2024a) and GenVidBench  (Ni et al., 2025) exhibit uncontrolled discrepancies in basic metadata—e.g., real videos are often longer than 10 seconds, whereas generated videos are typically under 4 seconds in GenVideo. Moreover, they reveal clear modality-level gaps in motion dynamics and content contrasts between real and generated videos. These differences introduce unintended shortcuts, enabling models to rely on superficial cues like duration or resolution rather than actual visual realism. As a result, VidGuard-R1 attains over 96% accuracy on both GenVideo and GenVidBench by exploiting such artifacts. To mitigate this reward hacking behavior, we construct a curated dataset with standardized video properties, encouraging models to focus on intrinsic visual content.
We collect real videos from the InternVid (Wang et al., 2024) and ActivityNet (Caba Heilbron et al., 2015) datasets and generate their corresponding fake counterparts using HunyuanVideo (Kong et al., 2024) and CogVideoX (Yang et al., 2024). We specifically choose these two models because they support conditioning on both the first-frame image and a text description—an essential requirement for generating videos that are contextually aligned with their real counterparts. To achieve such alignment, we provide the generation models with the first frame of each real video along with a textual caption describing its content. For ActivityNet, which lacks native captions, we extract concise descriptions using Qwen2.5-VL 72B. This pairing strategy mitigates content-based biases and forces the model to reason over subtle visual details.
@@ 3.3.3 GRPO with quality evolutionary videos (GRPO-Q)
This reward formulation enables the model to move beyond binary discrimination and perform fine-grained analysis of video quality. By learning to associate subtle differences in generation steps with quality variations, the model develops a deeper understanding of the diffusion process and its impact on perceptual realism. As a result, it can not only detect whether a video is fake, but also infer and estimate the degree of quality degradation in generated videos. This facilitates more interpretable and controllable evaluation of generated content quality.
@@ 4.1.3 Training Setup
We employ Qwen2.5-VL-7B as the base MLLM and conduct all experiments on four NVIDIA A100 GPUs (80GB). Each video is represented by up to 16 frames, each resized to $28\times 28$ and mapped to 128 feature channels for encoder input during both training and inference. For GenVideo and GenVidBench, we follow their official evaluation protocols and adopt 8-frame inputs. In GRPO training, we sample 8 responses per input; for GRPO-TA, we additionally sample 4 responses from temporally manipulated variants of the input to enhance robustness against temporal artifacts. Training proceeds in two stages: first, the base model is fine-tuned for one epoch on the CoT dataset, yielding the SFT-CoT MLLM; second, we initialize VidGuard-R1 with SFT-CoT and perform reinforcement learning for approximately 2,000 steps.
@@
VidGuard-R1 (GRPO, GenVideo-pretrained, Zero-shot) MLLM 97.24 96.59 99.88 99.93 88.14 96.37
@@ 4.2.1 Our dataset
We evaluate VidGuard-R1 on our dataset with several methods, including CNN-based models (SlowFast (Feichtenhofer et al., 2019), I3D (Carreira and Zisserman, 2017), TRN (Zhou et al., 2018)), Transformer-based models (UniFormer V2 Li et al. (2022a), TimeSformer (Bertasius et al., 2021), VideoSwin (Liu et al., 2022), MViT V2 (Li et al., 2022b)), and MLLM-based models (Qwen2.5-VL (Bai et al., 2025) and GPT-4.1 mini (OpenAI, 2025)). For CNN and Transformer models, we use the default training settings provided by the MMAction2 framework (Contributors, 2020).
@@ 4.2.2 GenVidBench benchmark
VidGuard-R1 (GRPO, GenVidBench-pretrained, Zero-shot) MLLM R 0.95 0.98 0.90 0.89 0.97 0.85 0.99 0.93 0.81 0.87 0.92
@@ 4.2.3 GenVideo benchmark
The GenVideo dataset comprises approximately 2.2M training samples and 20k testing samples, with generated videos sourced from a diverse set of models, including Sora (39), MorphStudio (35), Gen2 (Esser et al., 2023), HotShot (24), Lavie (Wang et al., 2025), Show-1 (Zhang et al., 2024a), MoonValley (34), Crafter (Chen et al., 2023), ModelScope (Wang et al., 2023a), and WildScrape (Wei et al., 2024). Following the official evaluation protocol, we benchmark two MLLM baselines and three variants of VidGuard-R1. Among these, VidGuard-R1 (GRPO) consistently outperforms almost all prior detection methods across videos generated by the various models. As shown in Table 3, it achieves an F1 score improvement of 0.06 compared to DeMamba-XCLIP. Complete detection model results are provided in Appendix B.
@@ 4.2.5 Ablation study
Method Expl. Score (HunyuanVideo) Expl. Score (CogVideoX)
@@
Explanation quality and accuracy comparison.
Table 4 presents results on the HunyuanVideo (Kong et al., 2024) and CogVideoX (Yang et al., 2024) datasets. We report explanation quality scores, which are rated on a 1–10 scale (with 10 indicating excellent quality and full alignment) by GPT-4.1 mini using the LLM-as-a-judge prompt described in Appendix D. Compared to baseline models such as Qwen2.5-VL-7B and GPT-4.1 mini, our VidGuard-R1 GRPO variants achieve consistent improvements in both classification accuracy and explanation quality.
@@
GRPO-TA reward ablation.
@@
Table 7: Cross-dataset evaluation for VidGuard-R1
Test Dataset Training Source Accuracy (%)
Ours VidGuard-R1 (Ours only) 81.65
Ours VidGuard-R1 (Ours + GenVideo) 82.97
GenVideo VidGuard-R1 (GenVideo only) 97.53
GenVideo VidGuard-R1 (Ours + GenVideo) 97.98
GRPO-Q reward ablation.
Table 7 presents an ablation study on GRPO-Q conducted on our dataset by varying the number of intermediate diffusion steps included per real video during fine-tuning. Using more steps provides richer supervision of video quality progression, improving detection accuracy. The best accuracy of 85.80% is obtained with five steps, which is the setting used in our main experiments.
Cross-dataset complementarity.
To assess whether training on a limited generative source induces overfitting, we conduct dataset-mixing experiments using VidGuard-R1 (GRPO). As shown in Table 7, augmenting our dataset with GenVideo leads to consistent performance gains across both evaluation sets, suggesting that the model benefits from heterogeneous training data and does not rely on artifacts from any single source. These findings indicate that incorporating diverse generative sources enhances overall accuracy, reinforcing that VidGuard-R1 learns generalizable detection cues rather than dataset-specific patterns.
4.3 Human Evaluation of Explanation Quality
To examine the coherence and interpretability of CoT rationales, we conducted a human evaluation of VidGuard-R1 (GRPO). Tables 9 and 9 summarize the two complementary studies.
Table 8: Label–rationale consistency and explanation quality for 20 videos
Metric Value
Annotators 5
Label–rationale alignment 89%
Rationale score (0–5) 3.9
Table 9: Human ranking of explanation quality
Model Avg. Rank
VidGuard-R1 (GRPO) 1.67
GPT-4o 2.08
Qwen2.5-VL-72B 2.22
Consistency and quality.
Five annotators evaluated twenty randomly selected fake videos that the model correctly identified. Annotators judged whether each rationale was consistent with the predicted label and assigned a quality score on a 0–5 scale after watching the corresponding video. As shown in Table 9, annotators reported 89% label–rationale agreement with an average quality score of 3.9. Lower scores ($\leq 2$) occurred primarily when the model emphasized subtle texture-level artifacts that were difficult for human raters to perceive.
Explanation quality ranking.
We conducted a blind ranking study in which five participants evaluated explanations from three models across twenty videos, consisting of ten real and ten synthetic videos. Table 9 shows that VidGuard-R1 achieved the strongest human preference, with the lowest average rank (1.67), outperforming GPT-4o and Qwen2.5-VL-72B. Participants consistently judged VidGuard-R1’s explanations to be more informative and higher quality. This improvement is likely because, although GPT-4o and Qwen2.5-VL-72B also describe video features well, their explanations are less effective at distinguishing real from synthetic content, whereas VidGuard-R1 better highlights the discriminative aspects relevant to this task.
Interpretation and broader implications.
Although the CoT rationales align closely with human judgments, they are not guaranteed to be correct in every case. Hallucinations may arise, and the rationales should therefore be shown with an appropriate disclaimer noting their AI-generated nature. Nevertheless, presenting interpretable cues helps users more effectively assess whether a video may be AI-generated, supporting a practical human-in-the-loop verification process.